# Floquet breathers in a modulated nonlinear lattice

### v2

@@ IV.D Stability of breathers in the time-modulated system
IV.E Extended autonomous symplectic system
The Hamiltonian structure of the dynamical equations, i.e., $\dot{u}_{n}=\partial H/\partial p_{n}$, $\dot{p}_{n}=-\partial H/\partial u_{n}$ is kept for the conjugate pairs of variables $u_{n}$, $p_{n}$, although the system is no longer autonomous. By defining $s=t$ or to be precise as (16) is $T_{b}$-periodic, as $s=\mathrm{rem}(t,T_{b})$, the remainder of $t/T_{b}$, i.e., $\mathrm{rem}(nT_{b}+t,T_{b})=t$ and $\mathrm{rem}(-nT_{b}-t,T_{b})=-t$, both for $n>0$ and $t>0$, so as $s$ is $T_{b}$ periodic. Then $\dot{s}=1$ as usual, and the system with the extra variable has become autonomous, but not symplectic as $s$ has no conjugate variable yet. The extra variable $h$ to make the system symplectic is obtained by defining a new Hamiltonian $K=H+h$, with $h$ the conjugate momentum of $s$ as $\dot{s}=\partial K/\partial h=1$ as previously obtained. Then, $\dot{h}=-\partial K/\partial s=-\partial H/\partial s=-\partial H/\partial t=-\textrm{d}H/\textrm{d}t$, or $h=E_{0}-H$, with $E_{0}$ the initial energy at t=0t=0Marthinsen and Owren (2016). The extended system is now symplectic and $\dot{u}_{b},\dot{s}]$ is now solution of (17) and being $T_{b}$ periodic, the Floquet multiplier +1+1 corresponding to the phase mode reappears. There is an extra +1+1 multiplier, because, symplectiness, implies that the multipliers are in pairs $\lambda,1/\lambda$, the extra mode corresponding to a solution with slightly different energy or growth mode.
The extended system reveals a hidden symplectic structure of the autonomous system with the implication that stability corresponds to all the Floquet multipliers being in the unit circle, but without the double multiplier of +1, which is recovered with the extended system. Figure 6-top-right shows these multipliers for $\Omega=2\omega_{\text{b}}$.
Note these conclusions are valid both the cases $\Omega>wb$ and $\Omega\leq\omega_{\text{b}}$ studied in the next two sections taking due care of the time of integration. They are also valid for space-time modulated systems with trivial changes that we will detail elsewhere.
V Breathers with time modulation and $\Omega=n_{T}\omega_{\text{b}}$
We consider the nonlinear dynamical system (9) with two main cases, $\omega_{\text{b}}=\Omega/n_{T}$ ($T_{\text{b}}=n_{T}T_{m}$), and $\omega_{\text{b}}=n_{T}\Omega$ ($T_{\text{b}}=T_{m}/n_{T}$), $n_{T}$ an integer. The first one corresponds to the well known phenomenon of period doubling, tripling, etc., while the latter correspond to a breather that oscillates several times faster than the modulation. According to the deduction in Sec. IV, fractional frequencies are also possible and we will consider them in Sec. VII.
There is no dissipation of energy, but as the system is parametrically forced the energy is conserved only at discrete times separated by the larger of the two periods, the breather period $T_{b}$ and the modulating one $T_{m}$. This is shown in Fig. 5 for $T_{b}=2T_{m}$.
The dynamical equations (9) are invariant under the transformation $t\rightarrow t+T_{m}$, therefore, if a phonon $(q,\omega)$, i.e., $u=\exp(\textrm{i}[qn-\omega t])$ is a solution of the linearized equation, then also $\exp(\textrm{i}m\Omega t)u$ is also a solution, or, in other words, there is another phonon band with frequencies $\omega^{\prime}=\omega+m\Omega$, with $m$ a positive or negative integer. But there are also two phonon bands, with positive and negative frequencies, then, if there is a phonon $(q,\omega)$, also there exists the phonon $(q,\Omega-\omega)$. As the unmodulated phonon band is $\in[\omega_{0}=1,\omega_{\text{max}}=1.01$, for $\Omega\simeq 2$,a new phonon band appears $\in[\Omega-\omega_{\text{max}}\simeq 0.98,\Omega-\omega_{0}\simeq 1]$. The phonon bands in the modulated system change position, and $\Omega=2\omega_{\text{b}}=1.96$, but those bands are very close with the possibility of interfering, depending on the exact value of $\omega_{\text{b}}$ and actual modulated phonons. This problem does not exist for $\Omega\geq 3$ as the band $\Omega-\omega_{\text{phon}}$ is much separated.
For $\Omega=2\omega_{\text{b}}$, and changing slightly $\omega_{\text{b}}$, there is an interval where the two phonon bands coincide and then the system becomes stimulated by $\Omega$ and the variables diverge. A dissipation term can be added, but as this is a very special case, we prefer not to modify the system and simply avoid those combination of frequencies.
V.A Period doubling $\Omega=2\omega_{\text{b}}$
The subject of time-modulation is of recent interest in photonic topological insulators. For example, Floquet solitons, for which the intensity repeats after each driving period, were found experimentally in optical waveguide arrays with Kerr nonlinearity Mukherjee and Rechtsman (2020) and period-doubled Floquet solitons were observed in a photonic topological insulator formed by an array of waveguides with a honeycomb pattern Mukherjee and Rechtsman (2023). For these solutions, the intensity period is doubled. The same happens in a helically modulated honeycomb lattice as well and a sinusoidally driven SSH lattice. Period-doubled solitons also appear in Pan and Wang (2020). These references show that the period-doubling phenomenon can appear in many different physical systems of interest and, in particular, photonic systems are ideal experimental devices to observe the connection between time modulation, topology and nonlinearity.
Period doubling also appears in our system. In figure 5-bottom we can see the evolution of the breather center and its neighboring sites, also the evolution of the energy, which is $T_{m}$-periodic and therefore also $T_{b}$-periodic. There is no net gain or loss of energy every period. The Floquet eigenvalues are also pictured in Fig. 6-top-left. The eigenvalues corresponding to the phonons form a pair of complex conjugate bands width frequencies $\omega_{n}=\omega_{\text{b}}+\theta_{n}/T_{b}$, as they have the same Krein signature Aubry (2006) they form a structurally stable subset. There is also a close but separate eigenvalue with a localized eigenvector shown in Fig. 6-bottom, and thus corresponding to the breather solution. It has the opposite Krein signature of the phonons, meaning it can collide with the phonon with the closest frequency leaving the unit circle, which indicates that the breather is unstable and another solution exists, most likely an extended one, corresponding to the merger of the breather with the upper phonon mode.
@@ V.A Period doubling $\Omega=2\omega_{\text{b}}$
V.B Cases $\omega_{\text{b}}=3\Omega$ to $\omega_{\text{b}}=10\Omega$
For $\omega_{\text{b}}=3\Omega$, the breather is slightly unstable with two real multipliers very close to +1. In spite of that, simulations can be done to simulate the breather evolution with relatively long time. The Floquet multipliers are represented in Fig. 7 and the Fourier spectrum in a system with noise and $N=160$ oscillators are shown.
For $\omega_{\text{b}}=n_{T}\Omega$, width $n_{T}=1\dots 10$, we have constructed breathers, for $n_{T}=3,4,8$ breathers are unstable, while the rest are stable, therefore there is no apparent pattern. For example, for $n_{T}=3$, there are two real eigenvalues slightly separated from +1+1, but, in spite of that, the breather has a long life.
VI Breathers with $\Omega=\omega_{\text{b}}/n_{T}$
A peculiarity of this case is that the stability of the breather and the Floquet matrix have to be calculated for a time equal to the modulation period $T_{m}=n_{T}T_{\text{b}}$ as the coordinates, and velocities have to repeat, but also the forces which depend on time and only repeat after $T_{m}$. Therefore, the phonon band appear with Floquet exponents multiplied by $n_{T}$ and spread out as they are given by $\theta_{q}=\omega_{q}T_{m}=n_{T}\omega_{q}T_{b}$ for a phonon of frequency $\omega_{q}$.
In this section we separate the study of the Sievers-Takeno mode Sievers and Takeno (1988) or single breather, from the Page mode Page (1990), or double breather. Values $n_{T}=1,2,\dots,32$ are considered and a breather obtained. For $n_{T}=32,\dots,40$ phonobreathers Marín and Aubry (1996) appear as explained in Sec. VI.B.
@@ VI.A Sievers-Takeno mode: single ILM with $\omega_{\text{b}}=n_{T}\Omega$
A stationary, single ILM or breather is a localized periodic vibration with a given frequency $\omega_{\text{b}}$ and with a single site, that we call the central one, having a large amplitude with respect to the phonons, and the amplitude of the neighbors diminishing rapidly with their distance to the center. They are often constructed from the anticontinuous limit with a single excited oscillator and the other at rest MacKay and Aubry (1994); Marín and Aubry (1996); Flach and Gorbach (2008)
@@
We construct the ILM in a small lattice with $N=16$ cantilevers, and check its stability by constructing the monodromy matrix and obtaining the Floquet multipliers and exponents Aubry (2006). Then, we embed the initial positions and within a larger lattice with $N=160$ particles with some small random noise velocities, let it evolve and perform the XTFFT to check the results. An example $\Omega=\omega_{\text{b}}/10$ with $\omega_{\text{b}}=0.98$ can be seen in Fig. 8. The top panel shows the profile at $t=0$ and the Floquet eigenvalues, showing its stability. The bottom panel illustrates the evolution of the variables for a period of the AC current.
We choose the case $\Omega=\omega_{\text{b}}/20$ with $\omega_{\text{b}}=0.98$, to illustrate the frequency-momenta representation as it is more easily seen. The top-right panel of Fig. 9-top-left shows the numerical XTFFT (see App. B.1) of an exact breather at zero temperature in a small system with $N=16$. The breather appears as a horizontal line corresponding to the breather frequency $\omega_{\text{b}}=0.98$, with two other breather bands at frequencies $\omega_{\text{b}}\pm\Omega$. The lines are extended to almost all the momenta because the breather is extremely localized. No phonons appear.
Figure 10: (Top) For $\Omega<\omega_{\text{b}}/33$, a phonobreather, appears with a background wave to the ILM. The inset shows the Floquet eigevalues with the Krein signature coded blue(+), red(-), black (zero). (Bottom) Fourier spectra. See text. Parameters: $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $\omega_{\text{b}}=0.98$.
Fig. 9-top-left shows the XTFFT of the same breather but located within a larger system with $N=160$ and thermalized with initial random velocities. The breather survives and the phonon band becomes apparent. Two replicas of the phonon band can be seen with frequencies $\omega=\pm\Omega+\sqrt{\omega_{0}^{2}+2\kappa(1-\cos(q)}$, and also two replicas of the breather band with frequencies $\omega_{\text{b}}\pm\Omega$. Each breather line replica is located below the corresponding phonon band replica. Other replicas with weaker intensity are out of the frame.
Fig.  9-bottom shows the numerical spectra obtained for the larger lattice with noise, adding up the intensities for all the particles at the same time. We can also see that also that the frequencies $\omega_{\text{b}}\pm\Omega$ appear displaced $\omega_{0}-\omega_{\text{b}}=1-0.98=0.02$ above and below the bottom of the secondary dispersion curves.
@@
Figure 12: (Top) ST in-phase breathers (Bottom) ST in-quadrature breathers. Evolution of the frequencies of phonons and breathers with indication of their stability. $\lambda_{\rm max}$ denotes the eigenvalue with the maximum absolute value. The breather/modulation frequency ratios correspond to $\omega_{\text{b}}/\Omega=m_{b}/m_{m}$ for $m_{b}=10$ and $m_{m}=[2,\dots,30]$, that is, $\omega_{\text{b}}/\Omega=m_{b}/m_{m}=[5,3.333,\dots,1,\dots,0.5,\dots,0.3333]$, Particular cases of interest are Floquet breathers $\omega_{\text{b}}=\Omega$ and period-doubled breathers $\omega_{\text{b}}=\Omega/2$. See text for details. Parameters: $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$.
@@ VI.B Phonobreathers: single ILM with phonon background for $\omega_{\text{b}}=n_{T}\Omega$
Although obtaining an ILM for increasing $n_{T}$ is possible, as $\Omega$ becomes smaller, the upper secondary ILM band $\omega_{\text{b}}+\Omega$ collides with the upper frequency of the phonon band $\omega_{\text{max}}=\sqrt{\omega_{0}+2\kappa(1-\cos(\pi))}=\sqrt{\omega_{0}+4\kappa}=\omega_{\text{b}}+\Omega$, or $\Omega=\omega_{\text{max}}-\omega_{\text{b}}$ and $\omega_{\text{b}}/\Omega\simeq 32.3$, for the values $\omega_{0}=1$ and $\kappa=0.0052$. Then for $n_{T}>30$ the upper phonons are excited and a extended wave with $q=\pi$ profile appears as the background of the ILM, with smaller but comparable amplitude, forming a phonobreather. A phonobreather is a periodic solution on a nonlinear system formed by a breather with a background of homogeneous amplitude. They were first described by Marín and Aubry Marín and Aubry (1996); Aubry (1997); Marín and Aubry (1998). The background is not a phonon as it is not a solution of the linearized system and has a significant amplitude, it can also be described as a nonlinear extended wave.
The profile of the phonobreather can be seen in Fig. 10-(Top) and at the bottom of the same figure, the Fourier spectra, where the colliding of the top of the phonon bands can be seen merging with the ILM bands of upper order, while in Fig. 9-bottom they are well separated.
@@ VI.C Page modes with $\omega_{\text{b}}=n_{T}\Omega$
A Page mode Page (1990); Flach and Gorbach (2008) is a double breather or ILM with two sites vibrating with the same amplitude, it can be bond-symmetric if both sites oscillate in-phase (iP), or bond-antisymmetric if they oscillate in anti-phase (aP). The iP mode is unstable and the aP is stable for an unmodulated Klein Gordon system with soft potential Archilla et al. (2003). Both two modes appear also in the modulated system with the same stability properties. Both mode profiles and Floquet multipliers are represented in Fig. 11 for $\omega_{\text{b}}=10\Omega$. The XTFFT of the in-phase mode shows the appearance of another breather band closer to the bottom of the phonon band, which correspond to a the unstable mode. Similar phenomena to the single ILM occur, i.e., formation of replicas of the phonon band and ILM band displaced $\pm\Omega$ in the frequency space and the excitation of the background forming a phonobr0eather when the secondary ILM band interacts with another replica of the phonon band just below.
@@
Figure 13: Floquet breathers: $\omega_{\text{b}}=\Omega$: Evolution of the moduli and angle of the Floquet multipliers with the breather frequency. The sign of the Krein signature of the eigenvalue is represented by the code: blue ++; red -; and black as zero. (Left) Breather in phase with the modulation; (Right) Breather in quadrature with the modulation. The in-phase breather (left) is unstable with the unstable eigenvalues in the positive real line. However, it can be continued until the two bands of phonon eigenvalues collide. The in-quadrature breather (right) is stable and structurally stable as the close eigenvalues have the same Krein signature and cannot abandon the unit circle. Eventually, the breather eigenvalue traverses the phonon bands with different Krein signatures bringing about oscillatory instabilities. It can be continued inside the phonon band. Parameters: $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $N=16$.
@@
Figure 14: Period-doubled breathers: $\omega_{\text{b}}=\Omega/2$: Evolution of the moduli and angle of the Floquet multipliers with the breather frequency. The sign of the Krein signature of the eigenvalue is represented by the code: blue ++; red -; and black as zero. (Left) Breather in phase with the modulation; (Right) Breather in quadrature with the modulation. The in-phase breather (left) is stable and structurally stable until its eigenvalue traverses the phonon band where oscillatory instabilities occur, and becomes again stable after going out of the phonon band. The crossing is apparent as the phonons have actually a $2\pi$ extra angle. It can be continued briefly until the two phonon bands collide bring about instabilites at +1+1 as the system itself becomes unstable due to the resonance between the positive and negative phonon bands. In spite of that the breather can be continued. The in-quadrature period-doubled breather (right) is unstable for all frequencies, with unstable multipliers in the positive real line. Parameters: $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $N=16$.

### v3

@@ V.A Period doubling $\Omega=2\omega_{\text{b}}$
V.B Cases $\omega_{\text{b}}=3\Omega$ to $\omega_{\text{b}}=10\Omega$
For $\omega_{\text{b}}=3\Omega$, the breather is slightly unstable with two real multipliers very close to +1. In spite of that, simulations can be done to simulate the breather evolution with relatively long time. The Floquet multipliers are represented in Fig. 7 and the Fourier spectrum in a system with noise and $N=160$ oscillators is also shown.
For $\omega_{\text{b}}=n_{T}\Omega$, width $n_{T}=1\dots 10$, we have constructed breathers, for $n_{T}=3,4,8$, breathers are unstable, while the rest are stable, therefore there is no apparent pattern. For example, for $n_{T}=3$, there are two real eigenvalues slightly separated from +1+1, but, in spite of that, the breather has a long life. In Sec. VII, there is also more information on the stability of this type of breathers and its evolution with the frequency.
VI Breathers with $\Omega=\omega_{\text{b}}/n_{T}$
A peculiarity of this case is that the linear stability of the breather and the Floquet matrix has to be calculated for a time equal to the modulation period $T_{m}=n_{T}T_{\text{b}}$ as the coordinates, and velocities have to repeat, but also the forces which depend on time and only repeat after $T_{m}$. Therefore, the phonon band appear with Floquet exponents multiplied by $n_{T}$ and spread out as they are given by $\theta_{q}=\omega_{q}T_{m}=n_{T}\omega_{q}T_{\text{b}}$ for a phonon of frequency $\omega_{q}$.
@@ VI Breathers with $\Omega=\omega_{\text{b}}/n_{T}$
In this section we separate the study of the Sievers-Takeno mode Sievers and Takeno (1988) or single breather, from the Page mode Page (1990) or double breather. Values $n_{T}=1,2,\dots,32$ are considered and breathers obtained. For $n_{T}=32,\dots,40$ phonobreathers Marín and Aubry (1996) appear as explained in Sec. VI.B.
@@ VI.A Sievers-Takeno mode: single ILM with $\omega_{\text{b}}=n_{T}\Omega$
A stationary, single ILM or breather is a localized periodic vibration with a given frequency $\omega_{\text{b}}$ and with a single site, that we call the central one, having a large amplitude with respect to the phonons, and the amplitude of the neighbors diminishing rapidly with their distance to the center. They are often constructed from the anticontinuous limit with a single excited oscillator and the others at rest.MacKay and Aubry (1994); Marín and Aubry (1996); Flach and Gorbach (2008)
@@
We construct the ILM in a small lattice with $N=16$ cantilevers, and check its linear stability by constructing the monodromy matrix and obtaining the Floquet multipliers and exponents.Aubry (2006) Then, we embed the initial positions within a larger lattice with $N=160$ particles with some small random noise velocities, let it evolve and perform the XTDFT to check the results. An example $\Omega=\omega_{\text{b}}/10$ with $\omega_{\text{b}}=0.98$ can be seen in Fig. 8. The top panel shows the profile at $t=0$ and the Floquet eigenvalues, showing its stability. The bottom panel illustrates the evolution of the variables for a period of the AC current.
We choose the case $\Omega=\omega_{\text{b}}/20$ with $\omega_{\text{b}}=0.98$, to illustrate the frequency-momenta representation as it is more easily seen. The top-left panel of Fig. 9 shows the numerical XTDFT (see App. C.1) of an exact breather at zero temperature in a small system with $N=16$. The breather appears as a horizontal line corresponding to the breather frequency $\omega_{\text{b}}=0.98$, with two other breather bands at frequencies $\omega_{\text{b}}\pm\Omega$. The lines are extended to almost all the momenta because the breather is extremely localized. No phonons appear.
Figure 10: (Top) For $\Omega<\omega_{\text{b}}/33$, a phonobreather appears with a background wave to the ILM. The inset shows the Floquet eigevalues with the Krein signature coded blue(+), red(-), black (zero). (Bottom) Fourier spectra. See text. Parameters: $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $\omega_{\text{b}}=0.98$.
Figure 9-top-right shows the XTDFT of the same breather but located within a larger system with $N=160$ and thermalized with initial random velocities. The breather survives and the phonon band becomes apparent. Two replicas of the phonon band can be seen with frequencies $\omega=\pm\Omega+\sqrt{\omega_{0}^{2}+2\kappa(1-\cos(q)}$, and also two replicas of the breather band with frequencies $\omega_{\text{b}}\pm\Omega$. Each breather-line replica is located below the corresponding phonon band replica. Other replicas with weaker intensity are out of the frame.
Figure 9-bottom shows the numerical Fourier spectra obtained for the larger lattice with noise, adding up the intensities for all the particles at the same time. We can also see that the frequencies $\omega_{\text{b}}\pm\Omega$ appear displaced $\omega_{0}-\omega_{\text{b}}=1-0.98=0.02$ above and below the bottom of the secondary dispersion curves.
@@ VI.B Phonobreathers: single ILM with phonon background for $\omega_{\text{b}}=n_{T}\Omega$
Although obtaining an ILM for increasing $n_{T}$ is possible, as $\Omega$ becomes smaller, the upper secondary ILM band $\omega_{\text{b}}+\Omega$ collides with the upper frequency of the phonon band $\omega_{\text{max}}=\sqrt{\omega_{0}^{2}+2\kappa(1-\cos(\pi))}=\sqrt{\omega_{0}^{2}+4\kappa}=\omega_{\text{b}}+\Omega$, or $\Omega=\omega_{\text{max}}-\omega_{\text{b}}$ and $\omega_{\text{b}}/\Omega\simeq 32.3$, for the values $\omega_{0}=1$ and $\kappa=0.0052$. Then for $n_{T}>30$ the upper phonons are excited and a extended wave with $q=\pi$ profile appears as the background of the ILM, with smaller but comparable amplitude, forming a phonobreather. A phonobreather is a periodic solution on a nonlinear system formed by a breather with a background of homogeneous amplitude. They were first described by Marín and Aubry.Marín and Aubry (1996); Aubry (1997); Marín and Aubry (1998) The background is not a phonon as it is not a solution of the linearized system and has a significant amplitude, it can also be described as a nonlinear extended wave.
The profile of the phonobreather can be seen in Fig. 10-top, and at the bottom of the same figure, the Fourier spectra, where the colliding of the top of the phonon bands can be seen merging with the ILM bands of upper order, while in Fig. 9-bottom they are well separated.
@@ VI.C Page modes with $\omega_{\text{b}}=n_{T}\Omega$
A Page mode Page (1990); Flach and Gorbach (2008) is a double breather or ILM with two sites vibrating with the same amplitude, it can be bond-symmetric if both sites oscillate in-phase (iP), or bond-antisymmetric if they oscillate in anti-phase (aP). The iP mode is unstable and the aP is stable for an unmodulated Klein Gordon system with soft potential.Marín et al. (1998) Both two modes appear also in the modulated system with the same stability properties. Both mode profiles and Floquet multipliers are represented in Fig. 11 for $\omega_{\text{b}}=10\Omega$. The XTDFT of the in-phase mode shows the appearance of another breather band closer to the bottom of the phonon band, which correspond to the unstable mode. Similar phenomena to the single ILM occur, i.e., formation of replicas of the phonon band and ILM band displaced $\pm\Omega$ in the frequency space and the excitation of the background forming a phonobreather when the secondary ILM band interacts with another replica of the phonon band just below.
Figure 12: (Top) ST in-phase breathers (Bottom) ST in-quadrature breathers. Evolution of the frequencies of phonons and breathers with indication of their stability. $\lambda_{\rm max}$ denotes the eigenvalue with the maximum absolute value. The breather/modulation frequency ratios correspond to $\omega_{\text{b}}/\Omega=m_{\text{b}}/m_{\text{m}}$ for $m_{\text{b}}=10$ and $m_{\text{m}}=[2,\dots,30]$, that is, $\omega_{\text{b}}/\Omega=m_{\text{b}}/m_{\text{m}}=[5,3.333,\dots,1,\dots,0.5,\dots,0.3333]$, corresponding to the slopes of the straight lines from left to right. Particular cases of interest are Floquet breathers with $\omega_{\text{b}}=\Omega$ and period-doubled breathers $\omega_{\text{b}}=\Omega/2$. See text for details. Parameters: $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$.
@@
Figure 13: Equal-period breathers: $\omega_{\text{b}}=\Omega$: Evolution of the moduli and angle of the Floquet multipliers with the breather frequency. The sign of the Krein signature of the eigenvalue is represented by the code: blue ++; red -; and black as zero. (Left) Breather in phase with the modulation; (Right) Breather in quadrature with the modulation. The in-phase breather (left) is unstable with the unstable eigenvalues in the positive real line. However, it can be continued until the two bands of phonon eigenvalues collide. The in-quadrature breather (right) is stable and structurally stable as the close eigenvalues have the same Krein signature and cannot abandon the unit circle. Eventually, the isolated eigenvalues corresponding to the internal modes of the breather traverse the phonon bands with different Krein signatures bringing about oscillatory instabilities. The breather can be continued inside the phonon band. Parameters: $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $N=16$.
@@
Figure 14: Period-doubled breathers: $\omega_{\text{b}}=\Omega/2$: Evolution of the moduli and angle of the Floquet multipliers with the breather frequency. The sign of the Krein signature of the eigenvalue is represented by the code: blue ++; red -; and black as zero. (Left) Breather in phase with the modulation; (Right) Breather in quadrature with the modulation. The in-phase breather (left) is stable and structurally stable until the isolated eigenvalues corresponding to internal modes of the breather traverse the phonon bands where oscillatory instabilities occur, and becomes again stable after going out of the phonon band. The crossing is apparent as the phonons have actually a $2\pi$ extra angle. It can be continued briefly until the two phonon bands collide bringing about instabilities at +1+1 as the system itself becomes unstable due to the resonance between the positive and negative phonon bands. In spite of that, the breather can be continued. The in-quadrature period-doubled breather (right) is unstable for all frequencies, with unstable multipliers in the positive real line. Parameters: $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $N=16$.